# DeblurDiNAT: A Lightweight and Effective Transformer for Image Deblurring

### v1

@@ Abstract
Blurry images may contain local and global non-uniform artifacts, which complicate the deblurring process and make it more challenging to achieve satisfactory results. Recently, Transformers generate improved deblurring outcomes than existing CNN architectures. However, the large model size and long inference time are still two bothersome issues which have not been fully explored. To this end, we propose DeblurDiNAT, a compact encoder-decoder Transformer which efficiently restores clean images from real-world blurry ones. We adopt an alternating dilation factor structure with the aim of global-local feature learning. Also, we observe that simply using self-attention layers in networks does not always produce good deblurred results. To solve this problem, we propose a channel modulation self-attention (CMSA) block, where a cross-channel learner (CCL) is utilized to capture channel relationships. In addition, we present a divide and multiply feed-forward network (DMFN) allowing fast feature propagation. Moreover, we design a lightweight gated feature fusion (LGFF) module, which performs controlled feature merging. Comprehensive experimental results show that the proposed model, named DeblurDiNAT, provides a favorable performance boost without introducing noticeable computational costs over the baseline, and achieves state-of-the-art (SOTA) performance on several image deblurring datasets. Compared to nearest competitors, our space-efficient and time-saving method demonstrates a stronger generalization ability with $3\%$-$68\%$ fewer parameters and produces deblurred images that are visually closer to the ground truth. The source code and pre-trained models are available at https://github.com/HanzhouLiu/DeblurDiNAT.git.
@@ 1 Introduction
Vision systems are widely applied to autonomous driving, medical research, and manufacturing [26]. It is essential to use high-quality sensing devices to capture clear visual data for them. Blurry photos are still unavoidable although camera modules develop fast [39]. Image deblurring is the procedure of removing blur artifacts from unclear or blurry images [4, 44], and is traditionally conceptualized as an inverse filtering problem [2, 12, 31]. Non-deep learning methods exhibit robust efficacy in certain scenarios [3, 1, 17, 43]. However, their performance is suboptimal in more intricate yet prevalent cases, for example, those involving strong motion blur [50].
@@
In the past decade, convolutional neural networks (CNN) play an important role in computer vision tasks [7, 49, 27]. The significant strength of CNNs is their ability to acquire generalizable image priors from large-scale datasets, making them become de facto for image deblurring [24, 38, 34, 5, 51, 21, 10, 29]. Nonetheless, the limited receptive field constrained by the kernel size and network depth prevents CNNs from modeling long-range pixel dependencies.
@@
Transformers, initially designed for natural language processing (NLP) [41], have adapted to handle visual data, challenging the dominance of CNNs in computer vision [9, 30, 46, 22]. Self-attention (SA), a fundamental component in Transformers, differs from the convolutional layer in CNNs and facilitates long-range feature learning through scaled dot-product. This attribute hinders the capacity of Transformers in capturing local blurred patterns, which frequently exist in real-word blurry images [13, 12, 40]. Also, the computational costs increase quadratically to process high-resolution images. To overcome these limitations, most methods divide image data into small spatial windows or strips. Although the model complexity is reduced and short-range blurring information is learnt, the deblurring performance may be negatively impacted due to the loss of spatial information. It has proven to be a challenging task for deblurring Transformers [42, 40, 47, 23] to achieve the optimal equilibrium between efficacy and efficiency.
@@
In this paper, we intend to find a balance that ensures optimized computation while maintaining high deblurring performance. For that purpose, we propose DeblurDiNAT, an effective and efficient deblurring Transformer, built on a novel hybrid encoder-decoder architecture. DeblurDiNAT produces high-quality deblurred images and shows a strong generalization ability, with compact size and fast inference time. Our SA strategy is based on Dilated Neighborhood Attention (DiNA) [14], a flexible and powerful sparse global attention pattern. DiNA introduces a dilation factor to allow controllable receptive fields without additional overhead [14, 15]. Similar to dilated convolutions [45], the dilation factor in DiNA indicates how much the kernel is widened, and a larger dilation factor results in a larger receptive field [20]. Through stacking Transformer blocks with small and large dilation factors alternately, our network learns both global and local blur information in an efficient way. To the best of our knowledge, this paper is the first work that comprehensively investigates the potential of DiNA in an image restoration task.
We propose a new SA scheme, channel modulation self-attention (CMSA), addressing a critical issue of applying SA to image data, i.e., SA lacking the ability to model cross-channel relationships effectively. CMSA uses element-wise multiplications to merge feature maps that are separately learnt by a dilated neighborhood attention block and a cross-channel learner (CCL). CCL draws inspiration from the channel attention method in SENet [19]. The significant difference is that rather than fully connecting all channels, CCL only considers each channel’s nearest neighbors, which is intuitively consistent with the rationale of neighborhood attention. We thoroughly compare our parallel design to variants with CCL at different positions.
@@
The other basic component of Transformers is the feed-forward network (FN). There are two common types of FNs in SOTA deblurring Transformers, FNs composed of fully connected layers followed by activation layers [40, 42], and the ones based on convolutional layers [47, 23]. In this paper, we aim to investigate and optimize the second type, where a non-linear GELU [16] function is used as part of gating mechanisms for controlled feature propagation. However, questions may arise regarding the effectiveness of GELU in this structure and none of them [47, 23] provide experimental data to analyze its impacts on deblurring performance. In this paper, we propose a divide and multiply feed-forward network (DMFN) without any non-linear activation function, thereby reducing computational loads. Experimental results show that this simple and clean design positively influences the overall deblurring outcomes.
Apart from the above modules, we visit a different element of the network, feature fusion, to further boost the deblurring power of our lightweight Transformer. We make the following empirical observations. Firstly, the conventional concatenate and convolve method may be a bottleneck of deblurring models. Secondly, existing SOTA feature fusion approaches [28, 6], originally designed for CNNs, add extra computational complexity to Transformers due to unneeded convolution, activation and normalization. Thirdly, in fusion blocks, positive impacts of SA are negligible. Motivated by the above observations, we present a lightweight gated feature fusion (LGFF) module for encoder-decoder Transformers, a universal method to merge multi-scale or same-scale features.
@@
  • We propose DeblurDiNAT, a compact Transformer for image deblurring, capable of capturing both global and local blurred patterns.
  • We present a channel modulation self-attention (CMSA) block, allowing the Transformer to learn complex multi-channel data in an efficient way.
  • We develop a divide and multiply feed-forward network (DMFN), which propagates features with optimal computational costs.
  • We design a lightweight gated feature fusion (LGFF) approach that is well-suited for both multi-scale and same-scale feature aggregation.
  • Extensive experiments show that DeblurDiNAT not only achieves SOTA performance, but also ensures efficiency and robustness.
@@ 2 Background
2.0.1 CNN-based Image Deblurring Models
Deep CNN-based methods have proven to be highly effective in image deblurring [18, 37, 33, 25, 38, 11, 48, 5, 32]. Tao et al. [38] investigated the coarse-to-fine scheme, and proposed a Scale-recurrent Network (SRN) with a simpler structure compared to conventional cascaded neural networks. Cho et al. [5] revisited the coarse-to-fine strategy and presented a multi-input multi-output UNet (MIMO-UNet) for fast image deblurring. Mehri et al. [32] chose to utilize multi-level representations and designed a multi-stage progressive architecture (MPRNet) to significantly enhance the deblurring performance.
2.0.2 Transformer-based Image Deblurring Models
Researchers have dedicated significant efforts to capturing spatial variant blur, in order to improve the applicability of Transformers in real-world image deblurring [42, 40, 47, 23]. Tsai et al. [40] leveraged intra-strip and inter-strip attention for estimating various blurred patterns and constructed Stripformer for dynamic scene image deblurring [40]. Kong et al. [23] advanced deblurring Transformers by introducing quantization and Fourier transform to self-attention and feed-forward networks. Although these approaches have produced better deblurring results than prevalent CNN models, they have not fully utilized cross-channel relationships in complex multi-channel data. Zamir et al. [47] proposed Restormer, an efficient Transformer which computes cross-covariance across feature channels rather than spatial dimension to model global connectivity. However, this architecture learns short-range dependencies exclusively through convolutional layers, which may fall short in effectively refining image details.
2.0.3 Multi-scale Feature Fusion
Li et al. [28] presented a Gated Fully Fusion (GFF) mechanism using multiple addition gates and element-wise multiplications to generate high-level feature maps for semantic segmentation. Dong et al. [8] proposed a Dense Feature Fusion (DFF) module for U-Net architectures, based on the back-projection technique. Nevertheless, redundant matrix operations in GFF and stacked convolutional layers in DFF result in heavy computational burdens. Cho et al. [5] adopted Asymmetric Feature Fusion (AFF) composed of only two convolutional layers to allow fast information flow from various scales across a U-Net. While the AFF design is simple, it is a limiting factor for deblurring models to learn complex representations in heavily blurred image data.
@@ 3 Method
Existing deblurring Transformers struggle with complexity and accuracy trade-off. To achieve SOTA performance with minimal space and time costs, we propose a Transformer-based architecture, DeblurDiNAT, as an effective, efficient and robust deblurring model. We firstly introduce DiNA [14] to image deblurring and present an alternating dilation factor structure to capture global and local blurred patterns. Then, we detail the novel channel modulation self-attention (CMSA), a basic component of DeblurDiNAT, leveraging the channel relationships for enhanced deblurring performance. After that, we present an easy while effective feature propagation method, divide and multiply feed-forward network (DMFN). Lastly, we provide a lightweight gated feature fusion (LGFF) module to further reduce computational costs and improve performance, which is suitable for both multi-scale and single-scale feature fusion. Fig. 1 demonstrates the overview of DeblurDiNAT architecture. Each encoder performs convolutions followed by Leaky ReLU activations, where convolutional layers capture spatial hierarchies in the input images and Leaky ReLU contributes to a robust learning process. Each decoder consists of cascaded Transformer blocks, in which global self-attention and local self-attention are adopted alternately to learn both long-range and short-range feature dependencies.
@@
$DiNA_{k}^{\delta}(i)=\emph{softmax}\>(\frac{\mathbf{A}_{i}^{(k,\delta)}}{\sqrt{d_{k}}})\mathbf{V}_{i}^{(k,\delta)}.$ (3)
Figure 1: Architecture of DeblurDiNAT for single image deblurring. DeblurDiNAT is a hybrid encoder-decoder network, where cascaded residual blocks extract multi-scale image features from input images, and alternating global and local Transformer blocks leverage hierarchical features for image reconstruction. Lightweight gated feature fusion (LGFF) modules perform effective and flexible feature fusion across the model.
@@ 3.1 Alternating Dilation Factor Structure
In DiNA, the lower bound for the dilation factor is $1$; the upper bound for the dilation factor is computed as $\lfloor\frac{n}{k}\rfloor$, the greatest integer less than or equal to $\frac{n}{k}$, where $n$ is the feature size, and $k$ is the kernel size. By incrementing the value of dilation factor, the self-attention window size is widened, which means longer-range dependencies are learnt by the network. To capture both local and global blurred patterns, we design a cascaded structure where the dilation factor $\delta\in\left\{1,\lfloor\frac{n}{k}\rfloor\right\}$ takes the lower bound value and upper bound value alternately. $k$ is set to a constant number $7$ in our design.
@@ 3.2 Channel Modulation Self Attention
Channel Modulation Self Attention (CMSA) is proposed as a powerful Transformer component which not only learns the spatial relationships between pixels but also captures channel dependencies. Fig. 2 (b)-(d) show that our proposed CMSA is composed of two units, DiNA and a cross-channel learner (CCL). CCL is essentially a lightweight channel attention module, which adaptively tunes the output feature maps of the self-attention layer according to the normalized input data. To optimize the extra runtime caused by CCL, we place it parallel to DiNA, and our extensive experiments show that this parallel design outperforms other variants, which are discussed in Sec. 4.3.3.
Given a layer normalized tensor $X\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$. The cross-channel learner (CCL) is computed as:
$CCL(X)=f^{-1}(Conv_{3}(f(AAP_{2d}^{1\times 1}(X)))),$ (4)
where $AAP_{2d}^{1\times 1}$ indicates a 2D adaptive average pooling, outputting a tensor of size $1\times 1$. $f$ is a function which squeezes and transposes a $C\times 1\times 1$ matrix, and results in a $1\times C$ matrix. $Conv_{3}$ denotes a 1D convolution with a kernel size of $3$. The batch dimension is ignored for simplicity in our discussions.
Then, the full process of CMSA is formulated as,
$\hat{X}=DiNA(X)\odot CCL(X),$ (5)
where $\odot$ denotes element-wise multiplication; $\hat{X}$ denotes the output tensor. $DiNA$ is illustrated in Eq. 1, Eq. 2, and Eq. 3.

### v2

@@ Abstract
While recent Transformer-based architectures generate improved deblurring outcomes than prior convolutional neural networks (CNNs), the large model size and heavy computational cost are still two bothersome issues which have not been fully resolved. Besides, the existing state-of-the-art (SOTA) approach trained on synthetic datasets may not perform well on a broad range of unseen real-world deblurring scenarios. To this end, we propose DeblurDiNAT, a compact and robust encoder-decoder Transformer which efficiently restores clean images from blurred ones. We adopt an alternating dilation factor structure to learn local and global blur patterns. Also, we observe that Transformers with conventional self-attention layers do not always produce good deblurred results. To solve this problem, we propose a channel modulation self-attention (CMSA) block, where a local cross-channel learner (LCCL) is utilized to capture channel relationships. In addition, we present a divide and multiply feed-forward network (DMFN) allowing fast feature propagation. Moreover, we design a lightweight dual-stage feature fusion (LDFF) module to introduce non-linearity. Comprehensive experimental results show that the proposed DeblurDiNAT provides a significant performance boost without a noticeable increase in computational cost over the baseline, and achieves SOTA performance on mainstream image deblurring datasets. Compared to nearest competitors, our space-efficient and time-saving model demonstrates a stronger generalization ability with $3\%$-$68\%$ fewer parameters and produces deblurred images that are visually closer to the ground truth. Source codes and pre-trained models are available at https://github.com/HanzhouLiu/DeblurDiNAT.git.
@@ 1 Introduction
Image deblurring is the procedure of removing blur artifacts from unclear or blurred images [5, 43], and is traditionally conceptualized as an inverse filtering problem [2, 13, 31]. Non-deep learning methods exhibit robust efficacy in certain scenarios [4, 1, 18, 42]. However, their performance is suboptimal in more intricate yet prevalent cases, for example, those involving strong motion blur [49].
@@
In the past decade, convolutional neural networks (CNN) play an important role in computer vision tasks [8, 48, 27]. The significant strength of CNNs is their ability to acquire generalizable image priors from large-scale datasets, making them become de facto for image deblurring [25, 37, 33, 6, 50, 22, 11, 29]. Nonetheless, the limited receptive field constrained by the kernel size and network depth prevents CNNs from modeling long-range pixel dependencies effectively.
@@
Transformers, initially designed for machine translation [39], have adapted to handle visual data, challenging the dominance of CNNs in computer vision [10, 30, 44, 23]. Self-attention (SA), a fundamental component in Transformers, differs from convolutional layers in CNNs and facilitates long-range feature learning through scaled dot-product. This attribute may hinder the capacity of Transformers in capturing local blurred patterns that exist in blurred images [14, 13, 38]. Besides, computational cost increases quadratically with respect to the input size. To overcome these limitations, most methods split image data into small spatial windows or introduce conventional 2-D convolutions [41, 38, 45, 24]. Although the model complexity is reduced and short-range visual representations are learnt in this way, the critical issue of partial loss of both spatial and cross-channel information has not yet been resolved. The existing SOTA deblurring model [24] tested on synthetic datasets is not able to generalize well from the known samples to a wide range of unseen real-world scenarios. It has proven to be a challenging task for deblurring Transformers to achieve the optimal balance among effectiveness, efficiency, and robustness.
In this paper, we intend to construct a deblurring Transformer with a compact size, low computational cost and SOTA generalization ability. We carefully investigate recent Transformer work for the deblurring task [41, 38, 45, 24] and propose three questions. First, is there an efficient way to capture both local and global blur patterns commonly existing in blurred images? Second, prior Transformers benefit from 2-D convolutions to learn cross-channel relationships globally [45, 24]. Is there an lightweight approach for Transformers to understand local channel dependencies that play an important role in image data either? Third, a simple gate without any gating mechanism functions effectively within a CNN architecture [3]. Does that hold true for a deblurring Transformer? If so, how to ensure non-linearity in order to model complex visual data? To answer these questions, we propose DeblurDiNAT, an effective, efficient and robust deblurring Transformer, built on a novel hybrid encoder-decoder architecture.
@@
As one of the core mechanisms of DeblurDiNAT, our SA strategy is based on Dilated Neighborhood Attention (DiNA) [15, 16], a flexible and powerful sparse global attention pattern. A dilation factor is used to adjust the receptive field size without additional overhead. A larger dilation factor results in a larger receptive field [21]. Through stacking Transformer blocks with small and large dilation factors alternately, our network learns both local and global blur information in an efficient way. To the best of our knowledge, this paper is the first work that comprehensively explores the potential of DiNA in an image restoration task.
To enable Transformers to understand local channel dependencies, we introduce a local cross-channel learner (LCCL) that only considers each channel’s nearest neighbors. LCCL draws inspiration from existing channel attention methods [20, 40]. Global channel relationships can be easily learnt by 2-D convolutions, which is not the focus of this paper. Formally, we propose a new SA framework, named channel modulation self-attention (CMSA), where a parallel LCCL dynamically modulates the outputs of an SA block (e.g., DiNA) based on its input feature maps. Although our parallel design was initially developed for efficiency, it outperforms other variants in terms of metric scores.
@@
The other basic component of Transformers is the feed-forward network (FN). There are two common types of FNs in SOTA deblurring Transformers, FNs composed of fully connected layers followed by activations [38, 41], and the ones based on 2-D convolutions [45, 24]. In this paper, we aim to investigate and optimize the second type, where a GELU [17] function is frequently used for the purpose of non-linearity. Questions may arise regarding the efficiency of this approach, considering that Transformers are built on sequentially stacked FNs and SA blocks. To further reduce computational loads, we present a divide and multiply feed-forward network (DMFN) without any non-linear activation function. Note that the non-linearity is handled in other modules. Experimental results show that this clean design not only decreases inference time but also positively influences deblurring outcomes.
Apart from the above novelties, we visit a different component of the network, feature fusion. We make the following empirical observations. First, existing Transformers for image deblurring only uses same-scale feature fusion, which cannot fully compensate for the information loss caused by the downsampling process, especially in deeper levels. Second, prior feature fusion approaches [28, 7], originally designed for CNNs, add extra computational complexity to Transformers due to unneeded convolutions, activations and normalizations. Third, as no activation exists in the decoder path of DeblurDiNAT, non-linearity should be introduced somewhere to handle and propagate complex visual data across the network. Therefore, we propose a non-linear lightweight dual-stage feature fusion (LDFF) module, an efficient method to merge multi-scale or same-scale features.
@@
  • We propose DeblurDiNAT, an encoder-decoder Transformer for image deblurring, capable of capturing both local and global blurred patterns.
  • We develop a channel modulation self-attention (CMSA) block, allowing the Transformer to learn local cross-channel relationships in an efficient way.
  • We present a divide and multiply feed-forward network (DMFN), which propagates features with optimal computational cost.
  • We design a lightweight dual-stage feature fusion (LDFF) approach that is well-suited for both multi-scale and same-scale feature aggregation.
  • Extensive experiments show that DeblurDiNAT is an efficient deblurring Transformer that performs on par with SOTA models, and demonstrate a very strong generalization ability on mainstream datasets.
@@ 2 Background
CNN-based Image Deblurring Models. Deep CNN-based methods have proven to be highly effective in image deblurring [19, 36, 32, 26, 37, 12, 47, 6, 46]. Tao et al. [37] investigated the coarse-to-fine scheme, and proposed a Scale-recurrent Network (SRN) with a simpler structure compared to conventional cascaded neural networks. Cho et al. [6] revisited the coarse-to-fine strategy and presented a multi-input multi-output UNet (MIMO-UNet) for fast image deblurring. Zamir et al. [46] chose to utilize multi-level representations and designed a multi-stage progressive architecture (MPRNet) to significantly enhance the deblurring performance.
Transformer-based Image Deblurring Models. Researchers have dedicated significant efforts to capturing spatial variant blur, in order to improve the applicability of Transformers in real-world image deblurring [41, 38, 45, 24]. Tsai et al. [38] leveraged intra-strip and inter-strip attention for estimating various blurred patterns and constructed Stripformer for dynamic scene deblurring [38]. Zamir et al. [45] proposed Restormer, an efficient Transformer which computes cross-covariance across feature channels rather than spatial dimension to model global connectivity. Although these approaches have produced better deblurring results than prevalent CNN models, they have not fully utilized cross-channel relationships in complex multi-channel data. Kong et al. [24] advanced deblurring Transformers by introducing quantization and Fourier transform to the basic building blocks, which achieves the highest metric score tested on the synthetic GoPro dataset [32]. However, it demonstrates a very weak generalization ability to process unknown real-world blurred images.
Multi-scale Feature Fusion. Li et al. [28] presented a Gated Fully Fusion (GFF) mechanism using multiple addition gates and element-wise multiplications to generate high-level feature maps for semantic segmentation. Dong et al. [9] proposed a Dense Feature Fusion (DFF) module for U-Net architectures, based on the back-projection technique. Nevertheless, redundant matrix operations in GFF and stacked convolutional layers in DFF result in heavy computational burdens. Cho et al. [6] adopted Asymmetric Feature Fusion (AFF) composed of only two convolutional layers to allow fast information flow across a U-Net, which might be a limiting factor to learn complex representations.
@@ 3 Method
DeblurDiNAT is an efficient hybrid Transformer for single image deblurring. In this section, we first introduce DiNA [15] to image deblurring and present an alternating dilation factor structure to capture local and global blurred patterns. Then, we detail the channel modulation self-attention (CMSA), leveraging local cross-channel relationships to model complex multi-channel data. After that, we present divide and multiply feed-forward network (DMFN) for fast feature propagation. Lastly, we provide a lightweight dual-stage feature fusion (LDFF) module to further reduce computational cost and improve image quality. Fig. 1 shows the architecture of DeblurDiNAT. Each encoder performs convolutions followed by Leaky ReLU activations, where convolutions extract hierarchical features and Leaky ReLU contributes to a robust learning process. The other components are explained in the following.
@@
$\mathrm{DiNA}_{k}^{\delta}(i)=\emph{softmax}\>(\frac{\mathbf{A}_{i}^{(k,\delta)}}{\sqrt{d_{k}}})\mathbf{V}_{i}^{(k,\delta)}.$ (3)
Figure 1: Architecture of DeblurDiNAT for single image deblurring. DeblurDiNAT is a hybrid encoder-decoder network, where cascaded residual blocks extract multi-scale image features from input images, and alternating local and global Transformer blocks leverage hierarchical features for image reconstruction. Lightweight dual-stage feature fusion (LDFF) modules perform effective and flexible feature fusion across the model.
@@ 3.1 Alternating Dilation Factor Structure
In DiNA, the lower bound for the dilation factor is $1$; the upper bound for the dilation factor is computed as $\lfloor\frac{n}{k}\rfloor$, the greatest integer less than or equal to $\frac{n}{k}$, where $n$ is the feature size, and $k$ is the kernel size. By incrementing the value of dilation factor, the self-attention window size is widened, which means longer-range dependencies are learnt by the network. To capture both local and global blurred patterns, we design a cascaded structure where the dilation factor $\delta\in\left\{1,\lfloor\frac{n}{k}\rfloor\right\}$ takes the lower bound value and upper bound value alternately. $k$ is set to a constant number $7$ in DeblurDiNAT.
@@ 3.2 Channel Modulation Self Attention
Fig. 2 (b) and (c) show that CMSA contains two parallel units, DiNA and a local cross-channel learner (LCCL). As illustrated in Fig. 2 (d), an LCCL first transforms the normalized 2-D features into 1-D data by global average pooling (GAP); then, it applies a 1-D convolution with a fixed kernel size to the intermediate features along the channel dimension; finally, a softmax function is adopted to compute attention scores. Tensor manipulations are ignored in Fig. 2 (d). Element-wise multiplications are used to merge the outputs of the LCCL and DiNA. We claim that our design has two benefits. First, as there is no channel reduction in the LCCL, information loss along the channel dimension is avoided. Second, the parallel architecture optimizes the extra runtime caused by the LCCL. Extensive experiments show that our CMSA outperforms other similar variants, which are discussed in Sec. 4.3.
Given an input tensor $\mathbf{X}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$. The output of CMSA is computed as,
$\hat{\mathbf{X}}$ $=\mathrm{DiNA}(\mathbf{X})\odot\mathrm{LCCL}(\mathbf{X}),$ (4)
${\rm LCCL}(\mathbf{X})$ $=f^{-1}(\mathrm{Conv}_{1d}(f(\mathrm{GAP}_{2d}^{1\times 1}(\mathcal{L}(\mathbf{X}))))),$
where $\odot$ denotes element-wise multiplication; $\mathrm{DiNA}$ is explained in Eq. 1, Eq. 2 and Eq. 3; $f$ is a tensor manipulation function which squeezes and transposes a $C\times 1\times 1$ matrix, resulting in a $1\times C$ matrix; $\mathrm{Conv}_{1d}$ denotes a 1D convolution with a kernel size of $3$; ${\rm GAP}_{2d}^{1\times 1}$ indicates global average pooling, outputting a tensor of size $1\times 1$; $\mathcal{L}$ is layer normalization.